OpenStudy (ohmyquaffle):

Line JK passes through points J(–4, –5) and K(–6, 3). If the equation of the line is written in slope-intercept form, y = mx + b, what is the value of b?

OpenStudy (phi):

I think we have to first find the slope m do you know how to find the slope between two points, using change in y divided by change in x ?

OpenStudy (ohmyquaffle):

I know I should but I do not, im not the most talented with math

OpenStudy (phi):

J(–4, –5) and K(–6, 3) do J-K for each number in the pair: -4 - (-6) and -5 - (3) can you simplify those ?

OpenStudy (ohmyquaffle):

2 and -8

OpenStudy (phi):

yes, the first number is the change in x i.e. 2 the 2nd number is the change in y -8 the slope m is change in y divided by change in x what do you get for m ?

OpenStudy (phi):

you should get m= -8/2 = -4 now in y = mx + b replace m with -4 y = -4 x + b to find b, use either (x,y) point for example (-4,-5). we replace x with -4 and the y with -5: -5 = -4*-4 + b -4 * -4 is +16 so that simplifies to -5 = 16+b the last step is to add -16 to both sides and then simplify.

OpenStudy (ohmyquaffle):

Thank you!